A group of students go out for lunch. If two have hamburgers and five have hot dogs, the bill will be $8.00. If five have hamburgers and two have hot dogs, the bill will be $9.50. What is the price of a hamburger?
Answer by mananth(16946)

A group of students go out for lunch. If two have hamburgers and five have hot dogs, the bill will be $8.00. If five have hamburgers and two have hot dogs, the bill will be $9.50. What is the price of a hamburger?
Let hamburgers cost $x each
Let hot dogs cost $y each
2x+5y==8.00
5x+2y=9.50
.
5(2x+5y)-2(5x+2y)= 5*8 - 2*9.50
10x+25y-10x -4y = 40-19
21y=21
y=1
.
plug value of y in any equation
2x+5*1 =8
2x=3
x=1.5
Hamburger cost is $ 1.5